RC meeting and Run Plan, Tuesday February 21th
From SBS wiki
Contents
RC Report
- https://logbooks.jlab.org/entry/4135876
- RC: Bogdan Wojtsekhowski, bogdanw@jlab.org
Summary
- Beam delivery is as usual
- We had ~45 uA beam some times
- Total charge collected over last three shifts is just 0.9C
Issues
- Average beam current is on the level of 1/3 of projected 45 uA
Current Status
- Production on He-3 target when beam is available
Run Plan
- Continue running 45 uA on He3
- Accelerator will prioritize beam and not have any maintenance days this week
Data Collection and Polarization Status
- 38.6C of uncorrected charge for this kinematic (44.0C of corrected charge which is 28% of the plan)
- Last NMR 51%
Other Topics/Status Updates
- 8 Target Operator Shifts need to be taken
- Readiness for Monday 3/27 target replacement
- Moller polarimeter readiness for 3/28 run
Requested Accesses to the Hall
- Holly need access related to SBS GEM electronics